# From the Stochastic Embedding Sufficiency Theorem to a Superspace Diffusion Framework

### 20 Mar 2026 v1

@@ B.11 Updated Summary: Comparison of the Two Formulations
Test C7 validates the $\rho_{\mathrm{KG}}$ formulation: the pipeline recovers $\hat{\sigma}/\sigma_{0}$ vs. $\gamma$ with power-law exponent $\alpha=-0.993\;[-1.018,\;-0.970]$ at 95% confidence, consistent with $\alpha=-1$ ($p=0.58$).
@@ B.12 Complete Logical Chain
The operational diffusion coefficient of a relativistic particle is suppressed by exactly $1/\gamma$ relative to its rest-frame value. This is not a convention: it is forced by the requirement that the Fokker–Planck equation reproduce the Lorentz-covariant Klein–Gordon probability current. Test C7 confirms $\hat{\alpha}=-0.993\pm 0.024$ (95% CI), consistent with the predicted exponent $\alpha=-1$. $\blacksquare$
@@ B.13 Status of Each Logical Step
Consistency with C7 ($\alpha=-1$) Validated $\hat{\alpha}=-0.993$ [1.018-1.018, 0.970-0.970] at 95% CI
@@ Appendix C Derivation of the Gravitational Diffusion Coefficient
This appendix derives the gravitational diffusion coefficient $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ from the equivalence principle, the quantum fluctuation-dissipation theorem, and dimensional analysis. The derivation proceeds in three stages: (i) the universal structure of the diffusion coefficient is extracted from nine validated physical domains (C1–C9), identifying $\sigma^{2}=\gamma\cdot\hbar\omega/m$ as the master formula; (ii) for massless fields, $\hbar$ and $\omega$ are shown to cancel from $\sigma$, leaving $\sigma=c\sqrt{\gamma}$ with $\gamma$ as the sole free parameter; (iii) gravity is shown to be the unique domain where dimensional analysis fixes $\gamma$ from fundamental constants alone, yielding $\gamma_{\mathrm{grav}}=1/t_{\mathrm{P}}$ and therefore $\sigma=c/\sqrt{t_{\mathrm{P}}}$. The result is equivalent to $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ for length-like variables, giving RMS metric fluctuations of exactly $\ell_{\mathrm{P}}$ per Planck time. In Planck units, $\tilde{\sigma}=1$ universally.
@@ C.3 Foundational Inputs
Of these three postulates, only Postulate C.3 introduces content not already established in standard physics. Proposition C.1 below shows that the self-bath structure follows from the Mori–Zwanzig projection formalism applied to the superspace Fokker–Planck equation; the full derivation and its application to galactic kinematic observables is given in a companion paper (Part II).
@@
The full derivation: including the explicit computation of $\gamma_{\mathrm{eff}}$ from the cubic vertex of the Einstein–Hilbert action and its identification with $1/t_{P}$ via the gravitational self-energy: is given in Part II, where it enters the galactic-scale predictions through the Mori–Zwanzig layer hierarchy. ∎
@@ C.4 The Structural Theorem
The fluctuation-dissipation theorem [5, 22] for a damped oscillator in contact with a bath at temperature $T$ gives $\sigma^{2}=2\gamma k_{B}T/m$ (the C2 result). At $T=0$, quantum mechanics replaces $k_{B}T$ with the zero-point energy $\tfrac{1}{2}\hbar\omega$. The factor of 2 combines with $\tfrac{1}{2}$ to give (C.86). This is confirmed numerically by C8 (QHO validation, $R^{2}=0.9999$, mean error $0.47\%$ across 9 frequency values). ∎
@@ C.6 The Main Theorem
Theorem C.6 (Gravitational Diffusion Coefficient).
@@ C.12 The Quantum Fluctuation-Dissipation Theorem
Writing this in velocity form $\mathrm{d}v=(-\omega^{2}x-\gamma v)\mathrm{d}t+\sigma\mathrm{d}W$ and matching the noise correlator $\langle\xi\xi^{\prime}\rangle=2m\gamma\hbar\omega\,\delta(t-t^{\prime})$ with the Itô convention $\langle\sigma^{2}\rangle=S_{F}/(2m^{2}\gamma)$ at the resonant frequency yields $\sigma^{2}=\gamma\hbar\omega/m$. C8 validates this result numerically to $0.47\%$.
@@ C.14.1 The Equivalence Principle and Gravitational Universality
In the physical domains C1–C9 (Observation C.1), the coupling rate $\gamma$ is a phenomenological parameter set by the environment: for example, the damping rate of a mechanical oscillator depends on the medium. In the gravitational setting this freedom disappears: the equivalence principle (Proposition C.15) requires gravity to couple universally to all energy-momentum, including its own, while dimensional closure (only $G$, $c$, $\hbar$ are available) forces $\gamma$ to be a universal function of $\omega$ constructible from Planck units alone. The derivation below shows that these two constraints determine $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ with no free parameters.
@@ C.15.5 Convergence of the Three Arguments
Point 3: independent confirmation from the perturbative mode sum. The Mori–Zwanzig memory kernel (Part II, Theorem A.2) uses the standard $|V|^{2}$ prescription and yields $\gamma_{\mathrm{eff}}\propto\omega^{3}$ mode by mode, confirming that the functional form is not an artefact of the linear prescription.

### 25 Mar 2026 v2

@@ B.11 Updated Summary: Comparison of the Two Formulations
Test C7 validates the $\rho_{\mathrm{KG}}$ formulation: the pipeline recovers $\hat{\sigma}/\sigma_{0}$ vs. $\gamma$ with power-law exponent $\alpha=-0.993$ with bootstrap 95% CI [0.999,0.988][-0.999,\;-0.988] from 50 independent simulations, consistent with $\alpha=-1$ at the $0.72\%$ level.
@@ B.12 Complete Logical Chain
The operational diffusion coefficient of a relativistic particle is suppressed by exactly $1/\gamma$ relative to its rest-frame value. This is not a convention: it is forced by the requirement that the Fokker–Planck equation reproduce the Lorentz-covariant Klein–Gordon probability current. Test C7 confirms $\hat{\alpha}=-0.993$ with bootstrap 95% CI [0.999,0.988][-0.999,\;-0.988] from 50 independent simulations, consistent with the predicted exponent $\alpha=-1$. $\blacksquare$
@@ B.13 Status of Each Logical Step
Consistency with C7 ($\alpha=-1$) Validated $\hat{\alpha}=-0.993$ [0.999-0.999, 0.988-0.988] at 95% bootstrap CI
@@ Appendix C Derivation of the Gravitational Diffusion Coefficient
This appendix derives the gravitational diffusion coefficient $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ from the equivalence principle and the quantum fluctuation-dissipation theorem. Dimensional analysis constrains the form of the result; the physical content is supplied by the FDT (validated by C8 to $0.27\%$) and the self-coupling mechanism (derived from the equivalence principle). The derivation proceeds in three stages: (i) the universal structure of the diffusion coefficient is extracted from nine validated physical domains (C1–C9), identifying $\sigma^{2}=\gamma\cdot\hbar\omega/m$ as the master formula; (ii) for massless fields, $\hbar$ and $\omega$ are shown to cancel from $\sigma$, leaving $\sigma=c\sqrt{\gamma}$ with $\gamma$ as the sole free parameter; (iii) gravity is shown to be the unique domain where dimensional analysis fixes $\gamma$ from fundamental constants alone, yielding $\gamma_{\mathrm{grav}}=1/t_{\mathrm{P}}$ and therefore $\sigma=c/\sqrt{t_{\mathrm{P}}}$. The result is equivalent to $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ for length-like variables, giving RMS metric fluctuations of exactly $\ell_{\mathrm{P}}$ per Planck time. In Planck units, $\tilde{\sigma}=1$ universally.
@@ C.3 Foundational Inputs
Of these three postulates, only Postulate C.3 introduces content not already established in standard physics. Proposition C.1 below shows that the self-bath structure follows from the Mori–Zwanzig projection formalism applied to the superspace Fokker–Planck equation; the full derivation and its application to galactic kinematic observables is the subject of subsequent work.
@@
The full derivation: including the explicit computation of $\gamma_{\mathrm{eff}}$ from the cubic vertex of the Einstein–Hilbert action and its identification with $1/t_{P}$ via the gravitational self-energy: is the subject of subsequent work. The identification $\gamma_{\mathrm{eff}}=1/t_{P}$ follows independently from dimensional uniqueness (Lemma C.5). ∎
@@ C.4 The Structural Theorem
The fluctuation-dissipation theorem [5, 22] for a damped oscillator in contact with a bath at temperature $T$ gives $\sigma^{2}=2\gamma k_{B}T/m$ (the C2 result). At $T=0$, quantum mechanics replaces $k_{B}T$ with the zero-point energy $\tfrac{1}{2}\hbar\omega$. The factor of 2 combines with $\tfrac{1}{2}$ to give (C.86). This is confirmed numerically by C8 (QHO validation, $R^{2}=0.9999$ across 9 frequency values, $0.27\%$ error from 50 independent runs at the reference frequency; Table 2). ∎
@@ C.6 The Main Theorem
Theorem C.6 (The Gravitational Diffusion Theorem).
@@ C.12 The Quantum Fluctuation-Dissipation Theorem
Writing this in velocity form $\mathrm{d}v=(-\omega^{2}x-\gamma v)\mathrm{d}t+\sigma\mathrm{d}W$ and matching the noise correlator $\langle\xi\xi^{\prime}\rangle=2m\gamma\hbar\omega\,\delta(t-t^{\prime})$ with the Itô convention $\langle\sigma^{2}\rangle=S_{F}/(2m^{2}\gamma)$ at the resonant frequency yields $\sigma^{2}=\gamma\hbar\omega/m$. C8 validates this result numerically to $0.27\%$ (Table 2).
@@ C.14.1 The Equivalence Principle and Gravitational Universality
In the physical domains C1–C9 (Observation C.1), the coupling rate $\gamma$ is a phenomenological parameter set by the environment: for example, the damping rate of a mechanical oscillator depends on the medium. In the gravitational setting this freedom disappears: the equivalence principle (Proposition C.15) requires gravity to couple universally to all energy-momentum, including its own, while dimensional closure (only $G$, $c$, $\hbar$ are available) forces $\gamma$ to be a universal function of $\omega$ constructible from Planck units alone. The derivation below shows that these two constraints determine $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ uniquely.
@@ C.15.5 Convergence of the Three Arguments
Point 3: independent confirmation from the perturbative mode sum. A perturbative computation of the Mori–Zwanzig memory kernel using the standard $|V|^{2}$ prescription yields $\gamma_{\mathrm{eff}}\propto\omega^{3}$ mode by mode, confirming that the functional form is not an artefact of the linear prescription.